martai:

Jason and his friend went to a restaurant for dinner. The food that they ordered cost $25. A 7% sales tax was applied to the total. After the sales tax was applied, Jason paid a 15% tip on the total bill. How much did Jason pay at the restaurant? A. $30.50 B. $30.76 C. $28.75

martai:

@xxdeppressionxx help

xxDeppressionxx:

The answer is simple just divide 15 and 7 then add it to 25. You get 28

xxDeppressionxx:

Plus 75 which is 28.75

martai:

OKI

martai:

THX U

xxDeppressionxx:

Np

xxDeppressionxx:

The math is \[15÷7=?\] then you get \[3~and~add~that~\to~25~which~is~28~plus~tax~which~is~28.75\]

Vocaloid:

Politely disagree with the previous response. The bill is 25, and there’s a 7% sales tax on the bill. 25 + (0.07 * 25) = 26.75 After the sales tax is applied, there’s a 15% tip on the new total. 26.75 + (0.15*26.75) = 30.76